Suppose we have a spaceship about the size of a typical ocean cruise ship today, which means it has a mass of about 140 million kilograms, and we want to accelerate the ship to a speed of 11 % of the speed of light. Suppose you want to generate the energy to get it to cruising speed using matter-antimatter annihilation.

How much antimatter would you need to produce and take on the ship?

Express your answer using two significant figures.

_____________kg
Answer #1

The mass of the ship is 1.4x108 kg.

The required velocity is 0.11c = 0.11x3x108 =  3.3x107 m/s

The energy required to bring the ship to this velocity is 1/2mv2 = 1/2x1.4x108 x(3.3x107)2 (Since this is the kintetic energy the ship should possess at the required speed).

Energy required = 8.1675x1022 J

Collision of matter and antimatter converts all the matter into energy using the formula E = mc2

Thus, mass of both matter and antimatter required will be m = E/c2 = 8.1675x1022/(3x108)2 = 9.1x105 kg

Only half of this is required to be antimatter. Thus, total antimatter required will be 910000/2 = 4.6x105 kg